# Increasing Profitability and Confidence by using Interpretable Model for Investment Decisions

### v1

@@ 4.2 Feature aggregation, Model forecasting and Trading
Figure 1: KSE-100 ten years closing index
Figure 2: EMA 55, 100 and 200 of KSE-100 index stocks
Fig. 1 presents the visualization of KSE-100 closing index of the cumulative stock assets over the 10 years of period with quarterly interval highlighting significant peaks and falls in the data.
@@
Figure 3: Potential trading signals from EMA-55 strategy
Figure 4: Potential trading signals from RSI strategy
Fig. 3 shows the trading signals generated overtime by our proposed model with the EMA-55 support/resistance strategy. It shows almost an overfitting trend and fluctuates rapidly with market dynamics and generates maximum buy sell signals, therefore, providing a resistance trend overall which is good for short investments and naive investors.
@@
4.3 Explanation for Daily Active Trading and -Short-term Trading Strategies
The aggregate statistical strategy approach achieves 63.36% of accuracy by employing the optimal weights on each technical indicator i.e., EMA200_Signal: 3, EMA55_Signal: 2, EMA100_Signal: 1, MACD_Signals: 1, RSI_Signals: 5 utilizing majority weighted voting system. Fig. 4(a) shows a visual representation of the mean SHAP values for each feature. It shows significant impact of EMA55 followed by EMA100 and then RSI, in case of daily active trading or scalping. We can observe rare resistance and support from EMA 200 in this strategy. On the other hand, for short-term trading strategy, by employing the optimal weights i.e., EMA200_Signal: 2, EMA55_Signal: 1, EMA100_Signal: 2, MACD_Signals: 1, RSI_Signals: 2, an accuracy of 71.88% is achieved. Fig. 4(b) shows the significant impact of EMA100 followed by EMA55 and then RSI, with aggregate weekly statistical weighted vote. This clearly suggests that EMA100 plays an important role in this strategy.
@@ 4.3 Explanation for Daily Active Trading and -Short-term Trading Strategies
Fig. 5(a) and 5(b) show the block plots for daily active trading and short-term trading strategies respectively. These provide valuable insights to the traders by not only showing the influential features and polarity of a particular instance, but also reveals the interaction of indicators by showcasing overlapping features contributing to the predictions. Moreover, the interpretability of plots is verified, as the SHAP values are highly correlated with the outcomes achieved through statistical strategies for 1-day and 5-day trading.
Figure 5: Visualization of feature importance for the proposed 1-day and weekly trading respectively
Figure 6: Block plot visualization for proposed daily and short-term trading strategies respectively
Fig. 6(a) and 6(b) show the force plots, representing instance-wise visualization for the corresponding daily active trading and short-term trading strategies. In devising a final trading move, force plots indicate the magnitude and direction of the individual feature impact on the final prediction, aiding in easy interpretation to each type of trader. We can see a decision of ‘hold a share’ for an instance in Fig. 6(a), depicting EMA55 and EMA100 have higher contribution in the decision, when in contrast to other technical indicators. On contrary, Fig. 6(b) depicts a trading move of ‘buy a share’ with significant contribution of the two technical indicators.
@@
Figure 7: Force plot visualization for proposed short-term trading strategy instances
Figure 8: Decision plot indicating significant contributors in over decision pattern using machine learning model explanation
When comparing the random forest classifier to our proposed model, it is worth noting that both models exhibited similar significant factors and contributions, as depicted in figures 4(a) and 8 for daily active investment decision. However, despite this similarity, the random forest classifier achieved an overall accuracy of 53% for one day and 55% for five days buy/sell decision forecasting. This clearly indicates that our proposed model outperformed random forest classifier in daily active trading and short-term investment strategies.
Figure 9: Force plot indicating significant contributors using machine learning model explanation
@@ 4.4 Case Study of PSX Stock Trading
To validate the effectiveness of the proposed trading strategy, we have devised a case study to demonstrate the application of short-term trading strategy, using automated investments in managing a stock trader’s portfolio. The investment span is four years, starting from 2018. The initial capital investment for the portfolio is set at PKR 10,000. It is important to consider that an investor’s portfolio is practically diverse in nature including investments in other commodities like equites and bonds [15] for potential investments. When placing a ‘buy order’ as a strong signal from our strategy, the portfolio investment is utilized to buy the shares. This approach ensures that the investor capitalizes on favorable market conditions for maximizing potential gains. To showcase the results of automated trading in opting ‘sell order’, two approaches are implemented.
In the first conservative hold and wait approach, an investor with a stance of avoiding hasty sell decisions places a ‘sell order’ after a week’s hold of a share. In this way, the potential risks of losses are minimized and a portfolio of PKR 11,063 over a period of four months stock investment is yielded as depicted in Fig. 10 below.
Figure 10: Investment returns for a conservative portfolio
In contrast, the aggressive investment approach places a ‘sell order’ as soon as a particular threshold of profit (i.e., 2% or more on an original investment) is achieved. This approach depicts an optimistic investor who aims to capture gains quickly and capitalize on shorter-term market fluctuations. A better investment is yielded utilizing this approach with a portfolio return of amount PKR 11,165. Fig. 11 depicts the period of investments with returns over the same period as given in the first approach.
Figure 11: Investment returns for an aggressive portfolio
The case study demonstrates the efficacy of the proposed trading strategy in short-term stock investments. By implementing the automated trading approaches presented in this study, investors can optimize their investment decisions and potentially achieve a profit of 35% over the course of one year stock investment which is significant as compared to other investment commodities such as equities and bonds. These findings highlight the importance of utilizing data-driven strategies and automation in stock trading to maximize returns while effectively managing risks.
@@ 5 Conclusion and Future Work
In this research work, we investigated an interpretable investment decision-making model leveraging SHAP based explainability technique. Although the proposed statistical model provides a solid foundation for inference and yielded better results in terms of forecasting trading signals and enhanced interpretation, yet there are certain limitations inherent to such classical approaches in contrast to the state-of-the-art Machine Learning models. In real trading scenarios, market stakeholders are not merely dependent on endogenous market information. In fact, their decision making is also influenced by various factors, including sentiments built from news and social media, fear and greed ratios, and external economic factors such as foreign investments and exchange rates like the Dollar-to-Pakistani-Rupee rate. On the other hand, Machine Learning models, although less interpretable in contrast, can identify complex patterns and can handle data heterogeneity more flexibly. Leveraging a post hoc interpretability technique like SHAP, therefore, provides a more robust solution for ML-based black box models. As a future work, we aim to employ diverse alternative data to bridge the gap between academic financial models and practical solutions. For this, a potential area of research is Natural Language Financial Forecasting (NLFF) that aims to process and interpret unstructured financial text data to generate meaningful forecasts and assist in investment decision-making [29]. It combines the power of language understanding with financial analytics to provide valuable insights and predictions in the financial domain. Another possible research direction for a deeper understanding of the market dynamics within specific sectors is to identify key drivers of stock price movements at a granular level. This could involve investigating the influence of various factors such as exogenous factors, sector-specific news, company-specific events, market sentiment, and price action patterns of individual stocks. Finally, we aim to enhance the existing explainability capability of our model by incorporating decision rules into the explanation process. In this way, a more structured and interpretable format to understand the reasoning behind the predictions is possible.
Code Availability
The source code for the reproducibility of the experiments conducted in this study is available upon request.
Conflict of Interest
@@ Acknowledgements
This research is funded by the Higher Education Commission’s National Research Program for Universities (NRPU) under project number 20-15756 and titled ”Generating Plausible Counterfactual Explanations using Transformers in Natural Language-based Financial Forecasting (NLFF)”. We are deeply grateful for HEC-NRPU financial support, which allowed us to conduct the experiments and analysis presented in this paper. This support has facilitated collaboration and knowledge-sharing among our research team and Falki Capital (Pvt) Ltd. (Equities brokerage house).

### v2

@@ 4.2 Feature aggregation, model forecasting and trading
Figure 1:
Fig. 1 presents the visualization of KSE-100 closing index of the cumulative stock assets over the 10-year of period with quarterly intervals highlighting significant peaks and falls in the data.
Figure 2:
@@
Figure 3:
Fig. 3 shows the trading signals generated over time by our proposed model with the EMA-55 support/resistance strategy. It shows almost an overfitting trend fluctuates rapidly with market dynamics and generates maximum buy-sell signals, therefore, providing a resistance trend overall which is good for short investments and naive investors.
Figure 4:
@@
4.3 Explanation for daily active trading and short-term trading strategies
The aggregate statistical strategy approach achieves 63.36% accuracy by employing the optimal weights on each technical indicator i.e., EMA200_Signal: 3, EMA55_Signal: 2, EMA100_Signal: 1, MACD_Signals: 1, RSI_Signals: 5 utilizing majority weighted voting system. Fig. 5(a) shows a visual representation of the mean SHAP values for each feature. It shows a significant impact of EMA55 followed by EMA100 and then RSI, in case of daily active trading or scalping. We can observe rare resistance and support from EMA 200 in this strategy. On the other hand, for short-term trading strategy, by employing the optimal weights i.e., EMA200_Signal: 2, EMA55_Signal: 1, EMA100_Signal: 2, MACD_Signals: 1, RSI_Signals: 2, an accuracy of 71.88% is achieved. Fig. 5(b) shows the significant impact of EMA100 followed by EMA55 and then RSI, with aggregate weekly statistical weighted vote. This suggests that EMA100 plays an important role in this strategy.
Figure 5:
Figure 6:
@@ 4.3 Explanation for daily active trading and short-term trading strategies
Fig. 6(a) and 6(b) show the block plots for daily active trading and short-term trading strategies respectively. These provide valuable insights to the traders by not only showing the influential features and polarity of a particular instance but also revealing the interaction of indicators by showcasing overlapping features contributing to the predictions. Moreover, the interpretability of plots is verified, as the SHAP values are highly correlated with the outcomes achieved through statistical strategies for 1-day and 5-day trading.
Fig. 7(a) and 7(b) show the force plots, representing instance-wise visualization for the corresponding daily active trading and short-term trading strategies. In devising a final trading move, force plots indicate the magnitude and direction of the individual feature impact on the final prediction, aiding in easy interpretation for each type of trader. We can see a decision of ‘hold a share’ for an instance in Fig. 7(a), depicting EMA55 and EMA100 have higher contribution in the decision, when in contrast to other technical indicators. On the contrary, Fig. 7(b) depicts a trading move of ‘buy a share’ with a significant contribution of the two technical indicators.
When comparing the random forest classifier to our proposed model, it is worth noting that both models exhibited similar significant factors and contributions, as depicted in figures 5(a) and 8 for daily active investment decisions. However, despite this similarity, the random forest classifier achieved an overall accuracy of 53% for one day and 55% for five days buy-sell decision forecasting. This indicates that our proposed model outperformed the random forest classifier in daily active trading and short-term investment strategies.
@@
Figure 7:
Figure 8:
Figure 9:
@@ 4.4 Case study of PSX stock trading
To validate the effectiveness of the proposed trading strategy, we have devised a case study to demonstrate the application of a short-term trading strategy, using automated investments in managing a stock trader’s portfolio. The investment span is four years, starting from 2018. The initial capital investment for the portfolio is set at PKR 10,000. It is important to consider that an investor’s portfolio is practically diverse including investments in other commodities like equities and bonds [15] for potential investments. When placing a ‘buy order’ as a strong signal from our strategy, the portfolio investment is utilized to buy the shares. This approach ensures that the investor capitalizes on favourable market conditions to maximize potential gains. To showcase the results of automated trading in opting for ‘sell order’, two approaches are implemented.
In the first conservative hold-and-wait approach, an investor with a stance of avoiding hasty sell decisions places a ‘sell order’ after a week’s hold of a share. In this way, the potential risks of losses are minimized and a portfolio of PKR 11,063 throughout four months of stock investment is yielded as depicted in Fig. 10 below.
Figure 10:
In contrast, the aggressive investment approach places a ‘sell order’ as soon as a particular threshold of profit (i.e., 2% or more on an original investment) is achieved. This approach depicts an optimistic investor who aims to capture gains quickly and capitalize on shorter-term market fluctuations. A better investment is yielded by utilizing this approach with a portfolio return of amount PKR 11,165. Fig. 11 depicts the period of investments with returns over the same period as given in the first approach.
Figure 11:
The case study demonstrates the efficacy of the proposed trading strategy in short-term stock investments. By implementing the automated trading approaches presented in this study, investors can optimize their investment decisions and potentially achieve a profit of 35% throughout one year of stock investment which is significant as compared to other investment commodities such as equities and bonds. These findings highlight the importance of utilizing data-driven strategies and automation in stock trading to maximize returns while effectively managing risks.
@@ 5 Conclusion and Future Work
In this research work, we investigated an interpretable investment decision-making model leveraging the SHAP-based explainability technique. Although the proposed statistical model provides a solid foundation for inference and yields better results in terms of forecasting trading signals and enhanced interpretation, there are certain limitations inherent to such classical approaches in contrast to the state-of-the-art Machine Learning models. In real trading scenarios, market stakeholders are not merely dependent on endogenous market information. Their decision-making is also influenced by various factors, including sentiments built from news and social media, fear and greed ratios, and external economic factors such as foreign investments and exchange rates like the Dollar-to-Pakistani-Rupee rate. On the other hand, Machine Learning models, although less interpretable in contrast, can identify complex patterns and can handle data heterogeneity more flexibly. Leveraging a post hoc interpretability technique like SHAP, therefore, provides a more robust solution for ML-based black box models. In future work, we aim to employ diverse alternative data to bridge the gap between academic financial models and practical solutions. For this, a potential area of research is Natural Language Financial Forecasting (NLFF) which aims to process and interpret unstructured financial text data to generate meaningful forecasts and assist in investment decision-making [29]. It combines the power of language understanding with financial analytics to provide valuable insights and predictions in the financial domain. Another possible research direction for a deeper understanding of the market dynamics within specific sectors is to identify key drivers of stock price movements at a granular level. This could involve investigating the influence of various factors such as exogenous factors, sector-specific news, company-specific events, market sentiment, and price action patterns of individual stocks. Finally, we aim to enhance the existing explainability capability of our model by incorporating decision rules into the explanation process. In this way, a more structured and interpretable format to understand the reasoning behind the predictions is possible.
Abbreviations
SHAP, SHapley Additive exPlanations; AI, Artificial Intelligence; XAI, Explainable AI; PSX, Pakistan Stock Exchange; KSE, Karachi Stock Exchange; CPI, consumer price index; EMA, Exponential moving averages; MACD, Moving Average Convergence Divergence; RSI, Relative Strength Index; NLFF, Natural Language Financial Forecasting.
Declarations
Availability of data and material
The data will be available upon request. The GitHub repository https://github.com/sahar-arshad/PSX-Interpretability.git contains the implementation of the algorithms and models discussed in this paper, allowing for reproducibility and further exploration of the methods used in this study.
Competing interests
@@
Authors’ contributions
Sahar Arshad: Conceptualization, methodology design and implementation, along writing of the original draft.
Seemab Latif: Supervision, formal methodology analysis and critical review of the manuscript.
Ahmad Salman & Saadia Irfan: Contribution in revisions and final approval of the version to be published
Funding
This research is funded by the Higher Education Commission’s National Research Program for Universities (NRPU) under project number 20-15756 and title 'Generating Plausible Counterfactual Explanations using Transformers in Natural Language-based Financial Forecasting (NLFF)'.
@@ Acknowledgements
We are deeply grateful for HEC-NRPU financial support, which allowed us to conduct the experiments and analysis presented in this paper. This support has facilitated collaboration and knowledge-sharing among our research team and Falki Capital (Pvt) Ltd. (Equities brokerage house). The authors also acknowledge the invaluable support and resources provided by the CPInS Lab at SEECS-NUST, which has been instrumental in facilitating the research and publication of this work